# Contrastive Learning and Adversarial Disentanglement for Task-Oriented Semantic Communications

### v2

@@ VI Experimental Evaluations and Discussion
In this section, we present the experimental setup used to evaluate CLAD. We start by describing the datasets used in our experiments, followed by a discussion of the baseline methods, neural architectures, and the experimental setup. Finally, we present detailed evaluations and analysis of the results. [^1] The source code, models and results are available at https://github.com/OmarErak/CLAD.
@@ VI-A1 Datasets
The Colored MNIST and Colored FashionMNIST datasets are extensions of the standard MNIST [37] and FashionMNIST [38] datasets, each consisting of 60,000 28x28 grayscale images. In Colored MNIST, handwritten digits (0-9) are overlaid on colored backgrounds, while in Colored FashionMNIST, clothing items from 10 categories (e.g., T-shirts, coats, shoes) are similarly displayed on colored backgrounds. The introduction of background colors adds additional task-irrelevant information, creating a more challenging setup for the model to disentangle task-relevant features relevant to classifying digits or clothing items from background-related attributes. Furthermore, incorporating background color labels enables the evaluation of attribute inference attacks, where an adversary is trained to predict background color. This setup provides insight into the model’s ability to protect against such attacks while maintaining disentanglement between task-relevant and task-irrelevant features. To further evaluate the scalability and robustness of the proposed method on more complex and natural datasets, we also incorporate CIFAR-10 [krizhevsky2009learningmultiple]. CIFAR-10 consists of 60,000 32x32 color images in 10 classes, including airplanes, cars, birds, cats, and other natural objects. Compared to MNIST-based datasets, CIFAR-10 introduces significantly more visual variability and semantic richness, making it a more challenging benchmark.
@@ VI-A2 Neural Network Architectures
In this work, we employ a deep neural network (DNN) architecture for the classification tasks across all datasets. These architectures consists of convolutional and fully connected layers, with a latent dimension $d=64$. The architectures, detailed in Table I and Table III, serve as the backbone for the task-relevant encoder (transmitter) and the downstream classification task (receiver). The same architecture is used across all baselines to ensure a consistent and fair comparison of methods. To compute the IRI, for each method we train multiple reconstructors that aim to reconstruct the original input $\boldsymbol{x}$ from the latent representation $\hat{\boldsymbol{z}}$. The reconstructor architectures are described in Table II and Table IV, and it is used consistently across all methods for IRI computations.
@@ VI-A3 Channel Conditions
We evaluate the performance of CLAD compared to baseline methods using an AWGN channel model due to its widespread adoption and ease of use. Specifically, we consider training and testing the models at identical SNRs, ranging from -6 dB to 12 dB. This setting allows us to assess the robustness of each method across different noise levels in a controlled manner. For each SNR value, we train the models over multiple runs and average the results to mitigate any randomness introduced during training.
TABLE I: DNN Structure for the transmitter (encoder) and the receiver (classifier) used for Colored MNIST and FashionMNIST
@@
TABLE II: Architecture settings for the reconstructors used with Colored MNIST and FashionMNIST to evaluate IRI
@@
TABLE III: DNN structure for the transmitter (encoder) and the receiver (classifier) used for CIFAR-10
@@
TABLE IV: Architecture settings for the reconstructors used with CIFAR-10 to evaluate IRI
@@
VI-B Baselines
In our experiments, we compare the proposed method CLAD against two baselines: DeepJSCC [14] and Variational Information Bottleneck (VIB) [16, 13]. These methods provide a benchmark for task-oriented communication systems, helping to evaluate the effectiveness of our approach in disentangling task-relevant and task-irrelevant features, and improving predictive accuracy through contrastive learning.
VI-B1 DeepJSCC
VI-B2 Variational Information Bottleneck (VIB)
The variational information bottleneck (VIB) framework aims to compress the input $X$ into latent representation $Z$ while retaining sufficient information for predicting $Y$. VIB balances $I(Z;X)$ and $I(Z;Y)$ via hyperparameter $\beta$. In our experiments, we provide results for different values of $\beta$ to illustrate how varying the trade-off between compression and task relevance impacts task performance, IRI, and privacy. The work in [13] is based on VIB, and therefore, we reproduce their results as a baseline comparison.
By comparing our method to these two baselines, we demonstrate how CLAD improves task-relevant feature extraction, minimizes task-irrelevant information, achieves better performance in downstream classification tasks and provides better privacy.
@@ VI-C Evaluation Metrics
To evaluate the effectiveness of CLAD, we utilize three key metrics: task performance through classification accuracy, IRI, and attribute inference attack accuracy. These metrics provide insights into how well the method performs in terms of classification accuracy, information retention, and privacy awareness. Additionally, we assess all methods at different channel SNRs to measure performance under various dynamic transmission conditions.
@@ VI-C1 Task Performance (Accuracy)
VI-C2 Information Retenetion Index (IRI)
@@ VI-C3 Attribute Inference Attack
Given the encoded representation $\hat{\boldsymbol{z}}$, the adversary seeks to predict the background color of the image. If $\hat{\boldsymbol{z}}$ contains significant task-irrelevant information, the adversary will be able to classify the background color with high accuracy. To evaluate this, we train a background color classifier (with the same architecture as the task relevant classifier in Table I) on the encoded representation $\hat{\boldsymbol{z}}$ and report its classification accuracy. A higher accuracy in the attribute inference attack implies more task-irrelevant information is retained in $\hat{\boldsymbol{z}}$, indicating weaker privacy guarantees.
TABLE V: Evaluation of different methods on Colored MNIST dataset at SNR = 12 dB.
@@
VI-D Results and Analysis
In this subsection we thoroughly analyze and discuss the performance of the proposed method CLAD against the two baselines, DeepJSCC and VIB, on all three aforementioned datasets.
TABLE VI: Evaluation of different methods on Colored FashionMNIST dataset at SNR = 12 dB.
@@ VI-D Results and Analysis
TABLE VII: Evaluation of different methods on CIFAR10 dataset at SNR = 12 dB.
@@
VI-D1 Colored MNIST Results
Fig.  4(a) shows the accuracy as a function of the SNR for each method. Here, it is evident that CLAD outperforms all the baselines at different channel conditions. This is a strong indication that maximizing $I(\hat{Z};Y)$ through contrastive learning yields better feature extraction and classification performance, even with undesirable and high noise conditions. As seen in Table V, CLAD outperforms both DeepJSCC and VIB, achieving the highest classification accuracy of 98.42% at SNR = 12 dB. This represents a 0.46% improvement over DeepJSCC (97.96%) and a 0.52% improvement over VIB at $\beta=0.001$ (97.90%). The higher accuracy of CLAD can be attributed to its ability to learn a more disentangled and task-relevant latent space using contrastive learning and adversarial disentanglement.
From Table V, CLAD achieves a very low IRI value of 0.039, drastically outperforming DeepJSCC (0.608) and VIB at $\beta=0.001$ (0.1931). This substantial difference shows that CLAD is effective at disentangling and removing task-irrelevant information from its encoded representations. This is further supported by the significantly lower adversarial accuracy of CLAD compared to DeepJSCC and VIB at $\beta=0.001$. From Fig. LABEL:fig:colored_mnist_reconstructions, we can visually see that image reconstructions from CLAD excel in retaining task-relevant information and preserving the core structure of the digits while discarding irrelevant information such as background color and other style features such as thickness or slant in the digit. This contrasts with DeepJSCCC and VIB, which retain unnecessary information such as background color and thickness and slant of the digit.
Interestingly, from Table V, VIB at lower $\beta$ values (e.g., $\beta=0.01$ and $\beta=0.1$) achieves IRI values that are closer to CLAD. However, the major drawback of VIB is that, while it improves privacy, informativeness and minimality, it comes at the cost of accuracy, especially as $\beta$ increases, as shown in the table. For example, at $\beta=0.1$, VIB achieves even lower IRI, however its classification accuracy drops to 93.29%, which is significantly lower than both CLAD and DeepJSCC. This demonstrates that CLAD maintains task performance while better preserving accuracy compared to VIB.
@@ VI-D1 Colored MNIST Results
VI-D2 Colored FashionMNIST Results
For the more challenging Colored FashionMNIST dataset, it is evident from Fig.  4(b), that CLAD shows even greater improvements across all SNR levels, with performance gains of up to 3% at higher SNR levels compared to other baselines. This demonstrates that CLAD becomes even more useful and essential for more complex computer vision tasks. Additionally, we visualize the noisy feature vector from the Colored FashionMNIST classification task using a two-dimensional t-distributed stochastic neighbor embedding (t-SNE) [39] in Fig. 5. Since CLAD effectively captures task-relevant features through contrastive learning and discards task-irrelevant information, it produces better class separation compared to DeepJSCC, as observed in the clearer clustering of feature representations.
In Table VI CLAD outperforms DeepJSCC with a 2.46% improvement and VIB at $\beta=0.001$ with 2.44% increase in accuracy. The IRI and adversarial accuracy results for Colored FashionMNIST reflect similar trends as observed in the Colored MNIST dataset. CLAD achieves a considerably lower IRI and lower adversarial accuracy compared to DeepJSCC and VIB at $\beta=0.001$, further confirming that CLAD is better at disentangling task-irrelevant information and maintaining privacy awareness. Fig. LABEL:fig:fmnist_reconstructions, visually illustrates the varying levels of informativeness and minimality in the encoded representations produced by each method, with respect to the input. In this image CLAD effectively removes most style features such as color, texture, and fine details, focusing solely on the basic shapes necessary for classification. For example, the CLAD reconstructions of the shoe, trousers, and shirt are reduced to simple grayscale silhouettes, discarding the vibrant colors and textures seen in the original and DeepJSCC versions. DeepJSCC retains style features such as color and subtle details such as the texture of fabric, making the reconstructions more complex but less focused on task-relevant information. VIB also retains some color and texture but introduces significant blurring, partially losing both class-relevant and style features. In contrast, CLAD’s grayscale, simplified images prioritize the task-relevant information necessary for downstream classification.
Once again, from Table VI, VIB achieves smaller IRI values at smaller $\beta$ settings but suffers a drop in accuracy, and the IRI continues to decrease as $\beta$ increases at the expense of even lower accuracy. This demonstrates that while VIB can be tuned to prioritize privacy and improve the minimality of the feature vector, it often struggles to maintain high classification performance.
VI-D3 CIFAR-10 Results
For the more complex CIFAR-10 dataset, Fig. 4(c) illustrates that CLAD continues to outperform baseline methods across all SNR levels. The performance gap becomes particularly notable at higher SNRs, where CLAD achieves up to 2% improvement in classification accuracy over DeepJSCC and VIB.
As shown in Table VII, CLAD delivers the highest classification accuracy of 92.33% at 12 dB, outperforming DeepJSCC by 1.13% and VIB (with $\beta=0.001$) by 0.84%. Moreover, CLAD achieves a significantly lower IRI compared to DeepJSCC, demonstrating its ability to retain task-relevant features while discarding redundant information, even in a dataset with rich visual complexity and no explicit sensitive attributes. These results reinforce CLAD’s robustness in encoding informative yet compact representations for more challenging computer vision tasks, without compromising classification performance.
@@ VII Conclusions
In this work, we have proposed a new task-oriented communication system, CLAD, based on contrastive learning and adversarial disentanglement. The proposed framework effectively disentangles task-relevant and task-irrelevant features, leading to enhanced performance in downstream tasks while simultaneously improving privacy as demonstrated using attribute inference attacks. Additionally, by disentangling task-relevant and task-irrelevant features, we reduce number of information bits that need to be transmitted. By leveraging contrastive learning, we enforce the model to capture discriminative task-relevant features, while adversarial disentanglement ensures that task-irrelevant features are separated, minimizing the mutual information between task-relevant and irrelevant representations. Furthermore, we present an approach to compute and compare the information retention between the input and encoded feature vectors across different methods, offering a practical and effective metric for the community to evaluate different task-oriented communication systems, which has been often neglected. Our results demonstrate that CLAD outperforms different state of the art task-oriented communication systems in terms of task performance, privacy, informativeness and minimality of the encoded feature representation. Future work can explore extending CLAD to handle different tasks or more complex multimodal data, such as video or audio streams, where task relevance might shift dynamically.

### v3

@@ VI Experimental Evaluations and Discussion
In this section, we present the experimental setup used to evaluate CLAD. We use image classification as a representative task to illustrate the core concept of the proposed methods, developing an end-to-end learning framework that extracts low-dimensional, task-relevant, privacy-preserving, and channel-robust latent representations for trustworthy 6G-IoT applications. Importantly, the proposed framework is not limited to image classification alone. We start by describing the datasets used in our experiments, followed by a discussion of the baseline methods, neural architectures, and the experimental setup. Finally, we present detailed evaluations and analysis of the results. [^1] The source code, models and results are available at https://github.com/OmarErak/CLAD.
@@ VI-A1 Datasets
The Colored MNIST and Colored FashionMNIST datasets are extensions of the standard MNIST [39] and FashionMNIST [40] datasets, each consisting of 60,000 28x28 grayscale images. In Colored MNIST, handwritten digits (0-9) are overlaid on colored backgrounds, while in Colored FashionMNIST, clothing items from 10 categories (e.g., T-shirts, coats, shoes) are similarly displayed on colored backgrounds. The introduction of background colors adds additional task-irrelevant information, creating a more challenging setup for the model to disentangle task-relevant features relevant to classifying digits or clothing items from background-related attributes. Furthermore, incorporating background color labels enables the evaluation of attribute inference attacks, where an adversary is trained to predict background color. This setup provides insight into the model’s ability to protect against such attacks while maintaining disentanglement between task-relevant and task-irrelevant features. To further evaluate the scalability and robustness of the proposed method on more complex and natural datasets, we also incorporate CIFAR-10 [41]. CIFAR-10 consists of 60,000 32x32 color images in 10 classes, including airplanes, cars, birds, cats, and other natural objects. Compared to MNIST-based datasets, CIFAR-10 introduces significantly more visual variability and semantic richness, making it a more challenging benchmark.
@@ VI-A2 Neural Network Architectures
To simulate realistic 6G-IoT scenarios, we adopt deep neural network (DNN) architectures for both the task-relevant encoder at the transmitter and the downstream classifier at the receiver. These networks consist of convolutional and fully connected layers, structured around a latent dimension $d$. The encoder and classifier architectures, outlined in Table II and Table IV, are used consistently across all evaluated methods to ensure fair and reproducible comparisons. To compute the IRI, we additionally employ reconstructor networks that attempt to recover the original input $\boldsymbol{x}$ from the latent representation $\hat{\boldsymbol{z}}$. The architectures for these reconstructors are provided in Table III and Table V, and are applied uniformly across all methods during IRI evaluation.
@@ VI-A3 Channel Conditions
We evaluate the performance of CLAD compared to baseline methods using an AWGN channel model due to its widespread adoption. Specifically, we consider training and testing the models at identical SNRs, ranging from -6 dB to 12 dB. This setting allows us to assess the robustness of each method across different noise levels in a controlled manner. For each SNR value, we train the models over multiple runs and average the results to mitigate any randomness introduced during training. In the following experiments, we simulate a constrained wireless edge scenario by setting the channel bandwidth to 12.5kHz and the symbol rate to 9,600 baud, reflecting practical limitations in edge communication environments.
VI-B Baselines
In our experiments, we compare the proposed method CLAD against three baselines: DeepJSCC [14], VIB and[16, 13] and information bottleneck and adversarial learning (IBAL) [20]. These methods provide a benchmark for task-oriented communication systems, helping to evaluate the effectiveness of our approach in terms of privacy and downstream task performance.
VI-B1 DeepJSCC
VI-B2 Variational Information Bottleneck (VIB)
The VIB framework aims to compress the input $X$ into latent representation $Z$ while retaining sufficient information for predicting $Y$. VIB balances $I(Z;X)$ and $I(Z;Y)$ via hyperparameter $\beta$. In our experiments, we provide results for different values of $\beta$ to illustrate how varying the trade-off between compression and task relevance impacts task performance, IRI, and privacy. The Variational Feature Encoding (VFE) method in [13] is based on VIB, and therefore the VIB results presented are synonymous to VFE.
VI-B3 Information Bottleneck and Adversarial Learning (IBAL)
IBAL is a task-oriented semantic communication approach that modifies the traditional VIB objective by incorporating an additional distortion constraint. Specifically, IBAL optimizes a composite loss function that balances the original variational information bottleneck objective with an MSE term that is maximized, which encourages poor reconstructions and thereby enhances resistance to model inversion attacks.
By benchmarking against these three baselines, we show that CLAD more effectively extracts task-relevant features, suppresses task-irrelevant information, enhances downstream classification performance, and offers stronger privacy guarantees.
TABLE II: DNN Structure for the transmitter (encoder) and the receiver (classifier) used for Colored MNIST and FashionMNIST
@@ VI-B3 Information Bottleneck and Adversarial Learning (IBAL)
TABLE III: Architecture settings for the reconstructors used with Colored MNIST and FashionMNIST to evaluate IRI
@@
TABLE IV: DNN structure for the transmitter (encoder) and the receiver (classifier) used for CIFAR-10
@@
TABLE V: Architecture settings for the reconstructors used with CIFAR-10 to evaluate IRI
@@ VI-C Evaluation Metrics
To assess the effectiveness of CLAD, we employ four key evaluation metrics: classification accuracy for task performance, IRI, attribute inference attack accuracy, and model inversion attacks for privacy assessment. Furthermore, all methods are evaluated across a range of channel SNRs to examine their robustness under dynamic transmission conditions.
@@ VI-C1 Task Performance (Accuracy)
VI-C2 Information Retention Index (IRI)
@@
Figure 6: Rate-distortion curves for (a) the Colored MNIST dataset, (b) the Colored FashionMNIST dataset and (c) the CIFAR-10 dataset. SNR is set to 12dB.
@@ VI-C3 Attribute Inference Attack
Given the encoded representation $\hat{\boldsymbol{z}}$, the adversary seeks to predict the background color of the image. If $\hat{\boldsymbol{z}}$ contains significant task-irrelevant information, the adversary will be able to classify the background color with high accuracy. To evaluate this, we train a background color classifier (with the same architecture as the task relevant classifier in Table II) on the encoded representation $\hat{\boldsymbol{z}}$ and report its classification accuracy. A higher accuracy in the attribute inference attack implies more task-irrelevant information is retained in $\hat{\boldsymbol{z}}$, indicating weaker privacy guarantees.
VI-C4 Model Inversion Attack
To further evaluate privacy leakage, we consider model inversion attacks, which aim to reconstruct the original input data $\boldsymbol{x}$ from the encoded representation $\hat{\boldsymbol{z}}$. This type of attack simulates an adversary that gains access to the transmitted latent representation and trains a decoder to recover the input image. In our setup, we follow a black-box attack scenario, where the adversary does not have access to the encoder or its training data. Specifically, we split the dataset such that 4/5 of the data is used to train and test the encoder and classifier models, while the remaining 1/5 is reserved for training the adversary. The adversary learns a mapping from $\hat{\boldsymbol{z}}$ to the corresponding input image using only this held-out subset. A visually accurate reconstruction indicates that $\hat{\boldsymbol{z}}$ still encodes significant low-level input features, implying weaker privacy preservation. Similar to [20], we use SSIM as a performance measure for the model inversion attacks.
TABLE VI: Evaluation of different methods on Colored MNIST dataset at SNR = 12 dB, under a latency constraint of $t\leq 6.67$ ms.
@@ VI-C4 Model Inversion Attack
TABLE VII: Evaluation of different methods on Colored FashionMNIST dataset at SNR = 12 dB, under a latency constraint of $t\leq 6.67$ ms.
@@
TABLE VIII: Evaluation of different methods on CIFAR10 dataset at SNR = 12 dB, under a latency constraint of $t\leq 6.67$ ms.
@@
Figure 7: Privacy evaluation comparing CLAD, VIB, IBAL, and DeepJSCC across different settings on the Colored MNIST dataset.
@@
VI-D Results and Analysis
In this subsection we thoroughly analyze and discuss the performance of the proposed method CLAD against the three baselines, DeepJSCC, VIB, and IBAL on all three aforementioned datasets.
VI-D1 Task-Oriented Classification Performance
We begin by evaluating the classification accuracy of each method at a fixed SNR of 12 dB under a latency constraint of $t\leq 6.67$ ms. The latency constraint can also be seen as an encoded feature vector with a maximum of 64 dimension. Tables VI, VII, and VIII summarize the results for Colored MNIST, Colored FashionMNIST, and CIFAR-10, respectively. CLAD consistently achieves the highest accuracy across all datasets, outperforming DeepJSCC, VIB (across various $\beta$ values), and IBAL.
For instance, on Colored MNIST, CLAD achieves 98.42% accuracy, surpassing DeepJSCC (97.96%) and the best-performing VIB configuration (98.01%) with better privacy metrics. On Colred FashionMNIST, CLAD outperforms all other baselines by 1.50-3.00%. On the more complex CIFAR-10 dataset, CLAD achieves 92.33%, outperforming DeepJSCC by 1.12% and VIB ($\beta=0.001$) by 0.84%.
We further evaluate classification robustness under dynamic transmission scenarios. As shown in Fig. 5, CLAD maintains higher accuracy across all SNR levels from -6 dB to 12 dB. This robustness is particularly crucial in 6G-IoT environments, where channel conditions fluctuate and low-latency, on-device inference is essential.
VI-D2 Rate-Distortion Tradeoff under Latency Constraints
In real-time 6G-IoT systems, minimizing latency while maintaining high task performance is critical. Since communication latency in our setup is determined by the dimension of the transmitted feature vector $\hat{Z}$, which is fixed across methods, we assess how efficiently each method encodes task-relevant information within that constraint.
Fig. 6 shows the rate-distortion curves, where distortion corresponds to classification error and rate is reflected by the latency budget. Despite identical feature vector sizes, CLAD consistently achieves lower distortion. This indicates that CLAD produces more compact and informative semantic representations by focusing on preserving task-relevant features, as a result, CLAD achieves superior downstream accuracy under the same latency constraints, yielding a more favorable rate-distortion tradeoff. This makes it particularly well-suited for efficient, privacy-aware communication in latency-sensitive 6G-IoT applications.
VI-D3 Privacy Evaluation: Information Retention, Attribute Inference, and Model Inversion
Privacy is critical for trustworthy communication in 6G-IoT systems, particularly when transmitting semantically rich representations over shared or noisy channels. To assess this, we evaluate how well each method limits task-irrelevant information leakage under varying conditions and attacks.
IRI and Attribute Inference. From Tables VI , VII and VIII, CLAD achieves the lowest IRI and attribute inference accuracy across all methods and datasets. On Colored MNIST, CLAD reaches an IRI of 0.039 and adversarial accuracy of 19.83%, significantly outperforming DeepJSCC (IRI of 0.608, adversarial accuracy 79.16%) and VIB (ranging from 0.3931 to 0.0123 in IRI as $\beta$ increases). While higher $\beta$ values in VIB reduce information retention and adversarial success, they also degrade task accuracy dropping to 93.29% at $\beta=0.1$ and 11.36% at $\beta=1$. In contrast, CLAD maintains both privacy and classification accuracy (98.42%). IBAL provides a lower IRI (0.1762) and adversarial accuracy compared to the best performing VIB results, however CLAD outperforms it in both privacy and task performance.
Similarly, on Colored FashionMNIST, CLAD achieves an IRI of 0.0587 and attribute inference accuracy of 19.33%, outperforming VIB (0.2562 IRI and 47.00% inference accuracy at $\beta=0.001$) and DeepJSCC (0.5958 IRI, 79.52% attribute inference accuracy). IBAL shows a decent privacy tradeoff (IRI 0.1762, inference 29.86% on MNIST), but CLAD surpasses it in both privacy and task performance. On the more complex dataset, CIFAR10, a similar trend is seen, as CLAD provides the best task accuracy and lowest IRI compared to all other baselines.
Privacy vs. SNR and Latency. The impact of varying SNR and latency on privacy is illustrated in Fig. 7. In Fig. 7(a), attribute inference success increases with SNR for all methods, as better channel conditions enhance signal fidelity, but CLAD consistently maintains the lowest inference accuracy and model inversion SSIM. For instance, at 12 dB, DeepJSCC shows attribute inference accuracy above 75% and SSIM exceeding 0.6, while CLAD remains below attribute inference accuracy 25% and under 0.05 SSIM. IBAL and VIB offer intermediate privacy performance, but are sitll outperformed by CLAD. Fig. 7(b) presents a comparison of privacy across different latency levels at a fixed SNR of 12 dB. Across all latency settings, CLAD consistently achieves the lowest attribute inference accuracy and SSIM, indicating strong resistance to both attacks.
VI-D4 Visual Analysis and Qualitative Comparisons
In addition to numerical metrics, Fig. 9 presents reconstructed images under each method. CLAD effectively removes stylistic and task-irrelevant details (e.g., color, background texture), focusing on shape and structure relevant for classification. VIB and IBAL reconstructions are blurrier, while DeepJSCC retains vivid background and texture details, making it vulnerable to inference attacks. These visual insights further support the quantitative findings. Furthermore, Fig. 8 shows a 2D t-SNE visualization of the encoded features for the Colored FashionMNIST dataset. The embeddings produced by CLAD result in more compact and clearly separated class clusters compared to DeepJSCC, indicating better task relevance and intra-class consistency. In contrast, DeepJSCC features show more overlap and spread.
@@ VII Conclusions
We proposed CLAD, a task-oriented communication framework that combines contrastive learning and adversarial disentanglement to extract compact, task-relevant features while improving privacy. By reducing task-irrelevant information in the latent space, CLAD enhances downstream performance and strengthens resistance to attribute inference and model inversion attacks. We also introduced the IRI, a practical and reproducible metric for comparing information preservation across methods. Extensive experiments across multiple datasets and channel conditions demonstrate that CLAD outperforms state-of-the-art baselines in terms of task accuracy, privacy, and minimality of representation, making it a suitable solution for semantic communication in 6G-IoT systems. Future work may explore extending CLAD to more complex input modalities and dynamically adaptive scenarios where task relevance can shift over time.